OpenStudy (anonymous):

Add the following: 4/3x+15 + 2x/x^2-25

OpenStudy (anonymous):

@Loujoelou :p

OpenStudy (anonymous):

Yay @agentx5

OpenStudy (anonymous):

Makalyla, can you put parenthesis around what you intended please? 2x/x^2-25 for example means, by order of operations \[\frac{2x}{x^2}-25\] But 2x/(x^2-25) is: \[\frac{2x}{x^2-25}\] Order of Operations (PELMDAS) 1. Parenthesis 2. Exponents & Logarithms 3. Multiplication & Division 4. Addition & Subtraction Make sense? Lots of people forget that on this website it seems, it would really help out all the people who volunteer to try to answer :-D

OpenStudy (anonymous):

What you're looking for to solve these is something called the Greatest Common Denominator, GCD for short.

OpenStudy (anonymous):

A:10x-20/3(x+5)(x^2-25) C: 8x/(3x+15)(x^2-25) B: 2x+4/3(x+5)(x-5) D:10(x-2)/3(x-5)(x+5)

OpenStudy (anonymous):

Assuming this: 4/(3x+15)+(2x)/(x^2-25) <-- parenthesis! Factor x^2 -25 to see it's (x+5)(x-5) GCD is then (3)(x+5)(x-5) so... \[\frac{10(x-2)}{3(x-5)(x+5)}\]

OpenStudy (anonymous):

So D. (:

OpenStudy (anonymous):

I see.